OpenStudy (anonymous):

Please Help 1. 6-8b=5b+30 2. 5y-2y=3y+2 3. 5x+2=2x-10 4. 4n- 8=3n+2 5. 1.2x+4.3=2.1-x 6.4.4s + 6.2 = 8.8s-1.8 7. 1/2b + 4= 1/8 +88 8. 3/4k-5=1/4-1 9. 8-5p=4p-1 10. 4b-8=10-2b 11. 0.2x-8=-2-x Please help

OpenStudy (anonymous):

bring all the variables to one side, bring all the consants to the other. divide by the coeficcient on the variable and arrive at the answer to each problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1. 6-8b=5b+30 -6 -6 -8b = 5b + 24 -5b -5b -13b = 24 /-13 /-13 b = 24/-13